A sample of 200 tourists in Florida shows 52 went to Disney, 57 went to Busch Gardens, and 52 visited both. Find the probability that a tourist attended neither Disney nor Busch Gardens.


1
Expert's answer
2021-08-09T15:46:10-0400
"P(D\cup BG)=P(D)+P(BG)-P(D\cap BG)"

"=\dfrac{52}{200}+\dfrac{57}{200}-\dfrac{52}{200}=0.285"




"1-P(D\cup BG)=1-0.285=0.715"

The probability that a tourist attended neither Disney nor Busch Gardens is "0.715."